OpenStudy (anonymous):

What is the inverse of f(x)=x^2-2x, x is greater than or equal to 1? Simplify the inverse and state the domain of f^-1?? and the inverse were its less than or equal to 1.

hero (hero):

@csmith3455

hero (hero):

To find the inverse of f(x), 1. set y = f(x), 2. swap variables x and y. 3. solve for y by completing the square 4. Replace y with \(f−1(x)\) 1. \(y=x2−2x\) 2. \(x=y2−2y\) 3. \(x+1=y2−2y+1\) \(x+1=(y−1)^2\) \(±\sqrt{x+1}=y−1\) \(±\sqrt{x+1}+1=y\) 4. The inverse is not a function Technically, the inverse of f(x) is not a function so we should not use the notation f−1(x)